The manager of a specialty food store combined almonds that cost $4.50 per pound with walnuts that cost $2.50 per pound. How many pounds of each were used to make a 100 pound mixture that costs $3.24 per pound?
Answer by richwmiller(17219)

a+w=100
4.5a+2.5w=100*3.24
a=37 w=63